OpenStudy (anonymous):

What is the sum of the solutions of 2[x - 1] - 4 = -2 ? A. 1 1/2 B. -2 C. 2 D. 0

OpenStudy (jiteshmeghwal9):

2(x-1)-4=-2 2x-2-4=-2 2x-6=-2 2x=6+2 2x=8 x=4

OpenStudy (anonymous):

@jiteshmeghwal9 is the answer B

OpenStudy (jiteshmeghwal9):

yes